OpenStudy (gabbyalicorn):

What is the value of the ratio of the surface area to the volume of the right prism? A. 1.9 B. 1.05 C. 0.9 D. 0.53

OpenStudy (gabbyalicorn):

 @gaos

OpenStudy (anonymous):

Look at the figure How many rectangle with 5 in and 4 in does it have?

OpenStudy (gabbyalicorn):

um 2?

OpenStudy (anonymous):

Yes, what's the surface area of each of them?

OpenStudy (gabbyalicorn):

4 x 2 base x hieght

OpenStudy (anonymous):

Base x height

OpenStudy (gabbyalicorn):

8

OpenStudy (anonymous):

Base=4 Height =5 What's the area?

OpenStudy (gabbyalicorn):

20

OpenStudy (anonymous):

Yes, each. But we have two rectangle, so the sum of the areas must be....

OpenStudy (gabbyalicorn):

4 and 5

OpenStudy (anonymous):

Two rectangle with an area of 20 each, 20+20=???

OpenStudy (gabbyalicorn):

40

OpenStudy (anonymous):

Yes, keep this number and let's calculate the second part

OpenStudy (gabbyalicorn):

k

OpenStudy (anonymous):

How many rectangle with 5x2 sides are there?

OpenStudy (gabbyalicorn):

10

OpenStudy (anonymous):

Yes and there are two of this . So 10+10=???

OpenStudy (gabbyalicorn):

20

OpenStudy (anonymous):

And the final part Two rectangle with 4x2 side What's the total area?

OpenStudy (gabbyalicorn):

8

OpenStudy (anonymous):

And there are two of these... So 8+8=???

OpenStudy (gabbyalicorn):

16

OpenStudy (anonymous):

Now, the final answer... The area of the prism is 16+40+20

OpenStudy (gabbyalicorn):

76

OpenStudy (anonymous):

Right, now let's calculate the volume Volume = width x height x base

OpenStudy (gabbyalicorn):

12,800

OpenStudy (anonymous):

No, V=5x4x2 V=???

OpenStudy (gabbyalicorn):

40

OpenStudy (anonymous):

Right, now let's calculate the percentage How much is 40/76=???

OpenStudy (gabbyalicorn):

0.526315

OpenStudy (gabbyalicorn):

D

OpenStudy (anonymous):

So the answer is letter....

OpenStudy (anonymous):

Perfect

OpenStudy (gabbyalicorn):

^^

OpenStudy (gabbyalicorn):

D

OpenStudy (gabbyalicorn):

Thank you!!!!!!!!!!!

OpenStudy (gabbyalicorn):

wait no it 1.9 ;-;